OpenStudy (acacia):

Determine if the function is even,odd, or neither. f(x) = x7 + x5

OpenStudy (ash2326):

If f(-x)= f(x), then function is even if f(-x)=-f(x), then function is odd. @acacia Can you find f(-x)?

OpenStudy (kirbykirby):

f(-x)=(-x)^7+(-x)^5 , now is f(-x) = -f(x) or f(x)

OpenStudy (acacia):

@ash2326 @kirbykirby so then would it be an even function?

OpenStudy (kirbykirby):

\(f(-x)=(-x)^7+(-x)^5=-x^7-x^5=-(x^7+x^5)\)

OpenStudy (kirbykirby):

do you recognize what's in the parentheses after the negative?

OpenStudy (acacia):

@kirbykirby yeah

OpenStudy (kirbykirby):

it's just f(x)! so we've shown that f(-x)=-f(x)